OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 15x - 5y = 30. intercept = (-2, 0); y-intercept = (0, 6) intercept = (0, -6); y-intercept = (2, 0) x-intercept = (2, 0); y-intercept = (0, -6) intercept = (0, 6); y-intercept = (-2, 0)

OpenStudy (zzr0ck3r):

to find the x intercept, let y = 0 and vice versa

OpenStudy (zzr0ck3r):

so if x is 0, what is y?

OpenStudy (anonymous):

15x-5y = 30 -5y = 30-15x y = 3x-6 y = 3(0)-6 y = 0-6 y = -6, (0,6) (0) = 3x-6 0 = 3x-6 -3x = -6 x = 2, (2,0) Solution: x-intercept = (2,0); y-intercept = (0,-6) Do you understand

OpenStudy (anonymous):

help with another?

OpenStudy (anonymous):

or a few lol

OpenStudy (anonymous):

sure

OpenStudy (zzr0ck3r):

you should try and figure it out, not just get the answers....

OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, 6) and has a slope of 4. y - 2 = 4(x + 6) y + 6 = 4(x - 2) y + 2 = 4(x - 6) y - 6 = 4(x + 2)

OpenStudy (zzr0ck3r):

no point in us helping if we are just giving answers.

OpenStudy (anonymous):

I showed you how to, ill help you but let me see what you know :D